OpenStudy (loser66):

I don't know how to start, Solve near point 0 ; x^2y'' -(x-2/9) y =0 Please, help

OpenStudy (anonymous):

\[x^2y''-\left(x-\frac{2}{9}\right)y=0\] or \[x^2y''-\frac{x-2}{9}y=0~~?\]

OpenStudy (loser66):

the first one

OpenStudy (tkhunny):

You WILL need a plan of attack. Series? Systems? Laplace Transform? What?

OpenStudy (loser66):

we don't have initial condition , how to use Laplace?

OpenStudy (anonymous):

@tkhunny, "near point x=0", so I would assume series. That, and the fact that OP has posted similar questions.

OpenStudy (anonymous):

It doesn't look like there are any non-zero solutions. At least, according to what I've worked through.

OpenStudy (loser66):

@SithsAndGiggles the problem doesn't have Q(x), or y' =0, how can I test whether 0 is singular regular point by taking limit?

OpenStudy (tkhunny):

@SithsAndGiggles I agree. Just trying to get @Loser66 to have an opinion and offer some reassurance that there is a chance of solving the problem.

OpenStudy (anonymous):

@Loser66, \(x=0\) could be an irregular singular point: https://en.wikipedia.org/wiki/Regular_singular_point#Examples_for_second_order_differential_equations

OpenStudy (loser66):

anyone, please

OpenStudy (loser66):

so far, I got, but don't know how to do next